### v2

# FastSLM: Hierarchical Frame Q-Former for Effective Speech Modality Adaptation

Abstract
Although Multimodal Large Language Models (MLLMs) have demonstrated remarkable capabilities in vision, language, and video understanding tasks, scaling them to long-form speech remains a critical bottleneck due to the explosive growth of input tokens. Existing speech-language models typically project high-frame-rate acoustic features directly into the LLM input space, rendering long-context processing computationally prohibitive as audio duration increases. In this paper, we present FastSLM, a token-efficient architecture designed to overcome this scalability limit through extreme temporal compression. At its core is the Hierarchical Frame Querying Transformer (HFQ-Former), which progressively distills local acoustic details into compact, semantically rich representations across multiple temporal scales. This hierarchical abstraction reduces the speech representation rate to just 1.67 tokens per second—achieving a 93% token reduction compared to standard frame-level adapters—while preserving the critical context required for complex reasoning. Experimental results demonstrate that FastSLM achieves competitive performance with state-of-the-art (SOTA) models on long-form benchmarks, despite operating with significantly lower FLOPs and parameter counts. Our findings establish that extreme token compression is a viable pathway to making real-time, long-context speech understanding feasible for LLMs, even under strict computational constraints. The source code and model checkpoints are available at https://anonymous.4open.science/r/FastSLM-8BD3.
Machine Learning, ICML
1 Introduction
Large Language Models (LLMs) (Achiam et al., 2023; Grattafiori et al., 2024; Comanici et al., 2025; Yang et al., 2025) have demonstrated remarkable reasoning capabilities, prompting a surge in research toward Multimodal LLMs (MLLMs) that perceive and understand the world through vision, audio, and video (Yin et al., 2024; Lyu et al., 2023). Among these modalities, speech is arguably the most primary and natural interface for human-AI interaction (Sakshi et al., 2024). Consequently, adapting LLMs to the speech domain has become a central focus, leading to the development of various Audio Language Models (ALMs) (Tang et al., 2024; Abouelenin et al., 2025; Chu et al., 2023, 2024; Goel et al., 2025).
However, a fundamental challenge hinders the scalability of current ALMs: the granularity mismatch between acoustic features and textual tokens. Prevalent approaches typically project frame-level features from encoders, such as Whisper (Radford et al., 2023) and Conformer (Gulati et al., 2020), directly into the LLM input space. While effective for short clips, this dense projection strategy generates hundreds of tokens for just a few seconds of audio. As the input duration extends to minutes or hours, the sequence length explodes, causing the computational cost of the autoregressive LLM, which scales quadratically ($O(N^{2})$), to become prohibitive. This phenomenon, often referred to as the token explosion, leads to excessive KV cache growth and unacceptable latency, rendering real-time processing impractical (Arif et al., 2025).
This bottleneck is particularly critical given the growing demand for long-form speech understanding in real-world applications such as meeting summarization (speech summarization (SSUM)) and spoken query-based question answering (SQQA) (Ghosh et al., 2024; Kang and Roy, 2024). Unlike short commands, these tasks require the model to reason over long temporal contexts containing thousands of acoustic events. Existing architectures, constrained by their dense token representation, struggle to handle such long sequences efficiently. To address this, we propose a paradigm shift from simple feature alignment to extreme temporal compression.
To bridge this gap, we present FastSLM, a cost-effective architecture designed to distill long-form speech into highly compact representations, making long-context reasoning feasible even under strict computational constraints. To drive this architecture, we introduce the Hierarchical Frame Querying Transformer (HFQ-Former), a novel mechanism that departs from flat projection by hierarchically compressing acoustic features. By progressively abstracting local details into global contexts, HFQ-Former reduces the speech representation rate to approximately 1.67 tokens per second, enabling the processing of hour-long audio on consumer-grade hardware. Furthermore, we introduce a three-stage training strategy designed to maximize data efficiency, enabling robust long-context adaptation without relying on expensive human-annotated long-form instruction datasets.
Our main contributions are summarized as follows:
  • We propose FastSLM, a streamlined architecture that solves the “token explosion” problem in MLLMs, enabling real-time reasoning over long-form speech with minimal computational overhead.
  • We introduce the HFQ-Former, which achieves extreme token compression (approximately 93% reduction vs. frame-level features) via hierarchical cross-attention, effectively balancing fine-grained acoustic perception with global semantic understanding.
  • We demonstrate a cost-effective three-stage training strategy that aligns pre-trained LLMs with long-form speech using only standard ASR corpora, bypassing the need for scarce long-context instruction data.
  • Empirical results show that FastSLM achieves competitive or superior performance to SOTA models on benchmarks such as SSUM and SQQA while exhibiting near-linear memory scaling.
2 Related Work
2.1 Audio and Speech Language Models
Although LLMs have achieved human-expert capabilities, their generation is inherently grounded in textual input. Consequently, to advance toward artificial general intelligence (AGI), numerous studies have explored extending LLMs to the audio (including speech) modality. Models such as AudioPaLM (Rubenstein et al., 2023), Kimi-Audio (Ding et al., 2025), Qwen-Audio (Chu et al., 2023), and Qwen2-Audio (Chu et al., 2024) demonstrate that frame-level speech features extracted from a speech encoder can be integrated into the LLM embedding space for end-to-end spoken understanding.
However, most of these models are primarily trained on short-form speech (typically under 30–60 seconds), limiting their ability to accurately process multi-minute inputs. To address this limitation, Voxtral-Mini, and Voxtral-Small (Liu et al., 2025a), Audio-Flamingo3 (Goel et al., 2025) introduced additional long-context training strategies and frame-level cross-attention for modality adaptation, enabling LLMs to better handle long-form speech. These works demonstrate that long-form speech reasoning is feasible when trained with sufficient data and computation.
Phi-4-Multimodal (Abouelenin et al., 2025), Gemini 2.5 (Comanici et al., 2025), and Qwen2.5-Omni (Xu et al., 2025) further show that LLMs can perform well across diverse modalities—including images, video, and speech—when trained on massive curated datasets. While powerful, their processing pipelines still depend on dense frame-level representations and incur substantial computational overhead for long-form speech inputs.
Notably, speech is one of the richest and semantically structured acoustic modalities. As a result, effective speech–LLM integration is essential for building systems capable of natural, interactive human communication. Despite strong progress, existing approaches typically rely on dense frame-level cross-attention or repeated window-level processing, both of which introduce substantial computational cost when handling multi-minute speech. Even models explicitly trained on long-form speech do not directly address the challenge of efficiently aligning long-form speech representations with the LLM under strict FLOPs constraints.
This gap motivates our work. Rather than merely enabling LLMs to process long-form speech, we aim to design a mechanism that efficiently compresses and aligns long-range speech features.
2.2 Q-Former for Speech Modality Adaptation
To align speech features extracted from a speech encoder with a LLM, several prior works have explored using a multi-layer perceptron (MLP) to project speech features into the LLM embedding space or inserting cross-attention layers directly into the LLM Transformer blocks. Although these approaches enable effective multimodal fusion, they require architectural modifications to the LLM and substantially increase parameter size and computational cost.
To overcome these limitations, recent studies have introduced Q-Former–based compression modules that convert high-frame-rate speech features into a small set of learnable query tokens. SALMONN (Tang et al., 2024) proposed a window-level Q-Former that compresses speech within fixed temporal windows by attending over frame-level features extracted from Whisper and CLAP encoder. Segment-level Q-Former (Yu et al., 2024) instead partitions the speech sequence into length-based segments, applies Q-Former compression to each segment, and concatenates the resulting tokens before feeding them to the LLM.
Beyond these early designs, MMCE-QFormer (Xue et al., 2024) introduced a multimodal context-enhanced Q-Former to jointly fuse speech and textual cues for decoder-based LLMs. CompressedToFindLM (Liu et al., 2025b) proposed a reference-guided compression strategy, where each compressed token is generated from local frame-level features using learned prototype representations. AlignFormer (Fan et al., 2025) addressed the temporal mismatch between speech and text by integrating a CTC layer with a dynamic-window Q-Former, providing better alignment for autoregressive decoding.
Existing Q-Former approaches (Li et al., 2023; Tang et al., 2024) typically utilize learnable queries to extract static features from fixed, short-term windows or simple segments (Yu et al., 2024). However, this design suffers from a context-granularity trade-off: applying Q-Formers to short windows risks missing global context, whereas applying them to long segments tends to dilute local acoustic details. Furthermore, naive segmentation leads to a linear increase in token count, failing to address the scalability challenges inherent in long-form speech. To resolve these issues, we propose the HFQ-Former, which introduces a progressive abstraction mechanism to efficiently capture both local details and global context.
3 Methodology
In Section 3.1, we describe the architecture of FastSLM and the inference process. In Section 3.2, we describe the training strategy employed for speech modality adaptation.
3.1 Model Architecture of FastSLM
FastSLM Overview: The overall architecture of FastSLM is illustrated in Fig. 1. Unlike conventional SLMs that directly concatenate dense acoustic features with text embeddings, FastSLM introduces a dedicated compression bottleneck to bridge the information density gap between modalities. The pipeline proceeds as follows: Raw waveforms are converted into Mel spectrograms (25-ms window and a 10-ms stride) and processed by a speech encoder (Radford et al., 2023) to extract high-fidelity frame-level features (50 Hz). Crucially, instead of feeding these dense frames to the LLM, we employ the HFQ-Former to distill them into highly compact semantic tokens. This mechanism achieves an extreme compression rate of approximately 1.67 tokens per second, reducing the sequence length by over 96% compared to the original encoder output. These compressed tokens are then concatenated with text prompts, allowing the LLM to perform long-context reasoning with minimal computational overhead.
Figure 1: Architecture of FastSLM.
Figure 2: The proposed flowchart of HFQ-Former. where, C denotes the concatenation.
HFQ-Former (Hierarchical Frame Querying Transformer): The HFQ-Former adopts a three-stage hierarchical design to progressively abstract speech information from acoustic to semantic levels. This mimics the cognitive process of progressively summarizing continuous acoustic signals into compact semantic representations. The detailed structure of HFQ-Former is shown in Fig. 2.
Stage 1 (Local Acoustic Extraction): The first stage focuses on preserving fine-grained local information. The high-frame-rate features $\hat{\mathbf{X}}_{1}^{A}\in\mathbb{R}^{F}$ from the encoder are fused with a set of learnable queries $\mathbf{Q}_{1}^{A}$ via cross-attention. Here, $\mathbf{Q}_{1}^{A}$ acts as a local probe, extracting phoneme-level details while maintaining the original temporal resolution. To preserve order during this dense interaction, we inject sinusoidal positional encoding $PE(\cdot)$ (He et al., 2022) into both queries and keys. The compression operation at stage $i$ is defined as:
$\text{Compressor}(\mathbf{Q}_{i}^{A},\hat{\mathbf{X}}_{i}^{A},\hat{\mathbf{X}}_{i}^{A})$ $=\text{Softmax}\!\left(\frac{\tilde{\mathbf{Q}}_{i}^{A}\tilde{\mathbf{X}}_{i}^{A\!T}}{\sqrt{d}}\right)$ (1)
$\cdot\hat{\mathbf{X}}_{i}^{A},\text{where }\tilde{\mathbf{Q}}_{i}^{A}$ $=\mathbf{Q}_{i}^{A}+PE(\mathbf{Q}_{i}^{A}),$
$\tilde{\mathbf{X}}_{i}^{A}$ $=\hat{\mathbf{X}}_{i}^{A}+PE(\hat{\mathbf{X}}_{i}^{A}),$
Stages 2 & 3 (Temporal Aggregation & Semantic Abstraction): Stage 1 captures local acoustic details, while Stages 2 and 3 progressively abstract this information into higher-level semantic representations. A Downsampler acts as a temporal filter, reducing redundancy (e.g., silence, stationary noise). Each Downsampler consists of dual convolutional layers with kernel size 3 and GELU activation (Hendrycks and Gimpel, 2016), applying a stride of 2 to halve temporal resolution at each step. Learnable queries $\mathbf{Q}_{i}^{A}$ at deeper stages attend to broader temporal contexts, transitioning from phoneme-level (Stage 1) to word/phrase-level (Stage 2), and finally to sentence-level semantics (Stage 3). The hierarchical feature update is formalized as:
$\begin{split}\hat{\mathbf{Q}}_{i}^{A}=\text{Compressor}_{i}(\mathbf{Q}_{i}^{A},\hat{\mathbf{X}}_{i}^{A},\hat{\mathbf{X}}_{i}^{A}),\\ \text{where}\quad\hat{\mathbf{X}}_{i}^{A}=\text{Downsampler}_{i}(\hat{\mathbf{X}}_{i-1}^{A}).\end{split}$ (2)
where $i\in\{2,3\}$, and $\hat{\mathbf{X}}_{0}^{A}$ denotes the initial encoder output. Multi-scale representations are concatenated as $\hat{\mathbf{Q}}^{A}=[\hat{\mathbf{Q}}_{1}^{A};\hat{\mathbf{Q}}_{2}^{A};\hat{\mathbf{Q}}_{3}^{A}]$ to form a comprehensive feature bank.
Semantic Distillation (Extreme Token Compression): Although $\hat{\mathbf{Q}}^{A}$ contains rich information, its length scales linearly with audio duration, which remains suboptimal for long-form reasoning. To achieve our target efficiency of 1.67 tokens/sec, we introduce a final Semantic Distillation step. Drawing inspiration from LLaVA-mini (Zhang et al., 2025), which compresses visual features into a single token, we hypothesize that speech segments can also be represented by a highly compact set of latent variables. However, unlike static images, speech represents a dynamic temporal stream with overlapping events. Therefore, instead of a single token, we employ a small, fixed number of learnable queries $\mathbf{Q}_{c}^{A}$ to distill the hierarchical information $\hat{\mathbf{Q}}^{A}$ into the final compressed tokens $\tilde{\mathbf{Q}}_{c}^{A}$:
$\tilde{\mathbf{Q}}_{c}^{A}=\text{C.Attn}\left({\mathbf{Q}}_{c}^{A},\hat{\mathbf{Q}}^{A},\hat{\mathbf{Q}}^{A}\right).$ (3)
Detail Recovery Mechanism: Extreme compression risks losing subtle but critical acoustic cues (e.g., intonation, speaker identity). To mitigate this, we incorporate a Detail Recovery attention block. This module allows the compressed semantic tokens $\tilde{\mathbf{Q}}_{c}^{A}$ to look back at the raw multi-scale source features $\hat{\mathbf{X}}^{A}$ and retrieve any missing fine-grained details before entering the LLM.
$\hat{\mathbf{Q}}_{c}^{A}=\text{Attention Block}\left(\tilde{\mathbf{Q}}_{c}^{A},\hat{\mathbf{X}}^{A}\right),$ (4)
where $\hat{\mathbf{X}}^{A}=[\hat{\mathbf{X}}_{0}^{A};\hat{\mathbf{X}}_{1}^{A};\hat{\mathbf{X}}_{2}^{A}]$. The final output $\hat{\mathbf{Q}}_{c}^{A}$ provides a strong balance between semantic density and acoustic fidelity, addressing the granularity mismatch described in Section 1.
To substantiate the efficacy of our hierarchical compression mechanism, we provide a qualitative analysis of cross-attention patterns in Appendix A. The visualizations demonstrate that HFQ-Former progressively shifts its attention toward deeper stages as the speech duration increases, confirming that the proposed hierarchical architecture is essential for effective long-range temporal abstraction. Furthermore, we present additional analyses on hierarchical temporal modeling in Appendix B and investigate the impact of the intermediate query size ($\hat{\mathbf{Q}}_{i}^{A}$) in Appendix C.
3.2 Three-Stage Training Strategy
To train FastSLM, we propose a three-stage speech modality adaptation strategy [Image of three-stage training pipeline for speech language models showing pre-training, long-form adaptation, and instruction tuning], designed to progressively enhance the model’s capability to understand and adapt to speech input. Across all stages, we adopt low-rank adaptation (LoRA) (Hu et al., 2022) to ensure cost-efficient training with minimal trainable parameters. Specifically, we set the LoRA rank to 16 and alpha to 64, resulting in a scaling factor of 4.
Pre-training (Short-form Speech Adaptation): In the first stage, the model is trained to adapt to short-form speech inputs. We construct a dataset of approximately 17K hours of speech-text pairs in both Korean and English, with each speech clip restricted to under 30 seconds. This ensures that the model can learn general ASR capabilities and effectively align speech with language. We adopt prompt formats inspired by hierarchical tags (Chu et al., 2023, 2024) to improve language-specific understanding during this stage. A detailed description can be found in Appendix D.
Long-form Speech Adaptation: Pre-trained speech encoders (Gulati et al., 2020; Radford et al., 2023; Elizalde et al., 2023; Chen et al., 2023) are typically limited to processing segments shorter than 30 seconds, hindering performance on long-speech tasks such as SSUM and SQQA. While Audio-Flamingo3 (Goel et al., 2025) addressed this by constructing bespoke instruction tuning datasets, such an approach is time-consuming and costly.
To provide a cost-effective alternative, we train the model on a curated ASR-based dataset containing speech–text pairs ranging from 1 to 15 minutes. This stage is designed not to train abstract reasoning directly but to strengthen the model’s fundamental ability to process extended speech sequences. Through long-form transcription training, the model learns to maintain temporal coherence and preserve acoustic features over lengthy contexts—a critical prerequisite for complex downstream tasks. The resulting long-context representations supply the LLM backbone with higher-quality, more coherent inputs, enabling superior performance on downstream tasks. Since ASR datasets are far more accessible than specialized instruction-tuning datasets, this approach offers a scalable path toward long-form speech modeling.
Instruction Tuning: In the final stage, we perform instruction tuning to enable the model to handle a variety of downstream tasks. Due to the scarcity of non-English multi-task datasets, we generated a Korean multi-task dataset using a text-to-speech (TTS) engine (Zhao et al., 2023), covering tasks such as SSUM and SQQA. Unlike previous stages, hierarchical language tags are no longer required as language identification is established; however, hierarchical task tags are employed to explicitly specify the task. Details are provided in Appendix D.
4 Experiment Results
4.1 Dataset Description
Pre-training Dataset: As described in Section 3.2, we constructed a bilingual dataset comprising 17K hours speech-text pairs to adapt ASR capabilities to the LLM during the pre-training stage. Including 9,152 hours of English speech-text pair (LibriSpeech (Panayotov et al., 2015), GigaSpeech-L (Chen et al., 2021), Voxpopuli (Wang et al., 2021), SpgiSpeech-M (O’Neill et al., 2021), Earnings-22 (Rio et al., 2022), AMI (Kraaij et al., 2005), Common Voice 15 (Ardila et al., 2019), AI-HUB ASR-En (The Open AI Dataset Project, 2021)) dataset, and 7,812 hours of Korean speech-text pair (AI-Hub-ASR-Ko (The Open AI Dataset Project, 2021)) dataset. A detailed description of the pre-training dataset can be found in Appendix E Table 7.
Long-form speech Dataset: To enhance the model capacity to process long-form speech input, particularly for tasks such as SSUM and SQQA, we constructed a dedicated long-form speech dataset. For English, we curated a total of 1,224 hours of long-form speech. For Korean, we curated a total of 1,012 hours of long-form speech.
Instruction Tuning Dataset: To enable robust instruction-following capabilities, we constructed a multi-task instruction tuning dataset covering four representative speech-language tasks: ASR, AST, SSUM, and SQQA. For ASR, we randomly sampled Korean and English speech-text pairs from the during pre-training. For SSUM, our dataset includes 1,600 hours of long-form dialogue from the MNSC corpus (Wang et al., 2025). As no public Korean SSUM dataset was available, we synthesized one by applying a TTS engine to the KMSS text summarization dataset (Kim et al., 2022). For SQQA, we used the English LibriSQA dataset (Zhao et al., 2024). For Korean, we constructed a parallel dataset by converting the text-based KorQuAD dataset (Lim et al., 2019) into speech via TTS. A detailed breakdown of datasets used for each instruction tuning task is provided in Table 1.
Table 1: Details of the instruction tuning dataset. “En” denotes English, “Ko” denotes Korean, and “En2Ko”, “Ko2En” indicate the translation directions.
Task Dataset Duration (hours) #Samples speech Language
ASR LibriSpeech GigaSpeech-S AI-HUB ASR 960 250 1500 281,241 230,068 320,000 En En Ko
AST AI-HUB AST (En2Ko) AI-HUB AST (Ko2En) 1,209 1,152 400,000 400,000 En Ko
SSUM SDS-PART6 KMSS 1,600 668 103,935 84,000 En Ko
SQQA LibriSQA KorQuAD-speech 364 483 104,014 100,243 En Ko
Total - 8,186 2,023,501 -
Evaluation Datasets: To evaluate the speech understanding capabilities of FastSLM, we conducted experiments across a variety of benchmark tasks.
  • ASR: For English, we used the OpenASR evaluation datasets (Srivastav et al., 2023). For Korean, we used the Common Voice 15 (Ardila et al., 2019) and Fleurs (Conneau et al., 2022) datasets, which are open datasets, for fair comparison of results. We evaluate transcription quality using character error rate (CER) for Korean and word error rate (WER) for English to reflect the linguistic characteristics of each language.
  • AST: We evaluated translation on the Fleurs (En2Ko/Ko2En) (Conneau et al., 2022) and Minds14 (Ko2En) (Gerz et al., 2021) datasets. We measured translation performance using the BLEU score (Post, 2018).
  • SSUM: Evaluation was conducted on SDS-PART6 (He et al., 2025) and KMSS-speech (Kim et al., 2022). Summarization quality was assessed using GPT-4 scoring with the LLM-as-a-judge framework (Zheng et al., 2023). Please refer to Appendix F for the SSUM judge prompt for GPT-4.
  • SQQA: We measured accuracy on the LibriSQA and KorQuAD-speech datasets to evaluate SQQA performance.
4.2 Experimental Setup
Model Architecture: FastSLM employs the encoder from Whisper-large-v3 (Radford et al., 2023) for speech feature extraction and adopts Qwen3-4B (Yang et al., 2025) as the backbone LLM for text generation. Despite its relatively compact size, Qwen3-4B exhibits sufficient capacity for comprehending speech-derived representations. In contrast to prior SLMs that typically utilize LLM backbone with 7 to 14 billion (B) parameters (Chu et al., 2023, 2024; Tang et al., 2024; Yu et al., 2024; Rubenstein et al., 2023; Ding et al., 2025; Liu et al., 2025a), FastSLM achieves a favorable cost-performance trade-off by leveraging lightweight architecture without compromising performance (Abouelenin et al., 2025; Ghosh et al., 2025). The HFQ-Former module within FastSLM compresses frame-level features via a hierarchical query-based mechanism. The number of learnable queries of $\mathbf{Q}_{i}^{A}$ was set to 80 cost-effectively (Yu et al., 2024), and the number of learnable queries of $\mathbf{Q}_{c}^{A}$ used as contextual input in LLM was set to 50 through the experiment. For further ablation studies and justifications for parameter selection, please refer to Section 4.5 Fig. 3.
[…]
The model implementation details and the training setup are summarized in Appendix H.
4.3 Comparison with Baseline Q-Former
To evaluate the performance of HFQ-Former, we compared it against two baseline methods: the segment-level Q-Former (SQ-Former) (Yu et al., 2024) and the window-level Q-Former (WQ-Former) (Tang et al., 2024), which explored speech modality adaptation using Q-Former. We additionally include an average pooling (AvgPool) baseline, in which frame-level speech features are downsampled through AvgPool and then projected into the LLM embedding space. This baseline allows us to evaluate whether direct downsampling can effectively replace a learnable compression module.
For evaluation, we employed the ASR task, as it provides a direct measure of how accurately the model can understand speech content. In addition, to assess the computational load imposed on the LLM when processing long-form speech, we measured the FLOPs of the LLM using a 5-minute speech input. All baselines were trained and evaluated under the same pre-training dataset, LLM backbone, LoRA configuration, and embedding dimensions to ensure a fully fair comparison. The detailed results are presented in Table 2.
Table 2: Comparison of WER across baseline methods. LS denotes the LibriSpeech.
Method Dataset LS-clean LS-other Voxpopuli #speech Tokens/Sec. LLM FLOPs (T)
AvgPool $\mathbf{1.88}$ $\mathbf{4.12}$ $7.14$ $25.0$ $30.6$
SQ-Former $2.32$ $4.87$ $8.37$ $2.67$ $3.32$
WQ-Former $2.14$ $4.51$ $7.26$ $2.93$ $3.65$
HFQ-Former (ours) $2.09$ $4.67$ $\mathbf{6.55}$ $\mathbf{1.67}$ $\mathbf{2.15}(-92\%)$
As shown in Table 2, HFQ-Former achieves the best WER on VoxPopuli and remains highly competitive on LS-clean, while using the fewest speech tokens per second (1.67 tokens/sec). Although AvgPool attains slightly lower WER on the LibriSpeech benchmarks, it requires a vastly larger number of speech tokens (25 tokens/sec) and incurs a substantially higher LLM computation cost (30.6 TFLOPs).
In contrast, HFQ-Former achieves a strong balance between accuracy and efficiency: it reduces the token rate by 38% compared to WQ-Former and lowers LLM FLOPs by 41.1% (3.65T → 2.15T), while still improving WER on VoxPopuli. These results indicate that HFQ-Former provides a significantly more efficient speech-to-LLM alignment mechanism, greatly reducing the computational burden of autoregressive decoding for long-form speech without compromising recognition quality.
Table 3: Comparison of FastSLM with other SLMs on various tasks. Bold indicates the best performance, and underline indicates the second-best performance. WER and CER are lower-is-better ($\downarrow$); Accuracy (ACC), BLEU, and Score are higher-is-better ($\uparrow$). N/A indicates the model lacks the corresponding capability. ‘*’ denotes results fine-tuned on an additional Korean dataset provided in the official Hugging Face supplementary material. Detailed ASR benchmark WER for each dataset is reported in Appendix I.
Task Metric Dataset FastSLM 4.8B Whisper 1.5B Qwen2-Audio 8B Phi4-MM 5.8B Voxtral-mini 4.7B Gemini-2.5- Flash
#speech tokens/30 Sec. 50 1500 101 375 750 960
ASR (En) WER $\downarrow$ OpenASR $\underline{6.47}$ $7.44$ $7.43$ $\mathbf{6.14}$ $7.05$ $9.29$
ASR (Ko) CER $\downarrow$ Fleurs Common Voice 15 $\mathbf{3.82}$ $7.92$ N/A N/A N/A $\underline{4.55}$
AST (En2Ko) BLEU $\uparrow$ Fleurs $\underline{7.39}$ N/A N/A *$2.62$ N/A $\mathbf{13.4}$
AST (Ko2En) BLEU $\uparrow$ Fleurs $\mathbf{19.5}$ $18.6$ N/A *$10.4$ N/A $\underline{19.2}$
AST (Ko2En) BLEU $\uparrow$ Minds14 $\underline{28.9}$ $\mathbf{29.5}$ N/A *$14.8$ N/A $26.3$
SSUM (En) Score (1-7) $\uparrow$ SDS-PART6 $5.40$ N/A $4.54$ $5.30$ $\underline{5.48}$ $\mathbf{5.87}$
SSUM (Ko) Score (1-7) $\uparrow$ KMSS $\underline{4.12}$ N/A N/A N/A N/A $\mathbf{4.37}$
SQQA (En) ACC $\uparrow$ LibriSQA $\mathbf{69.5}$ N/A $57.2$ $64.5$ $48.9$ $\underline{67.0}$
SQQA (Ko) ACC $\uparrow$ KorQuAD- speech $\mathbf{64.9}$ N/A N/A N/A N/A $\underline{64.8}$
4.4 Quantitative Results
We primarily compare FastSLM with strong speech-centric baselines such as WhisperV3 and AST, which directly align with our speech-only setting. For completeness, we additionally evaluate several multimodal models (Qwen2-Audio (Chu et al., 2024), Phi-4-Multimodal (MM) (Abouelenin et al., 2025), Gemini-2.5-Flash (Comanici et al., 2025), and Voxtral-Mini (Liu et al., 2025a)) in their speech-only mode. These models were not originally designed as SLMs, but we include them for an upper-bound comparison.
FastSLM demonstrates a powerful combination of efficiency and performance, achieving top-tier results with just 50 speech tokens per 30-second input—a fraction of that used by models like Whisper (1,500) and Gemini-2.5-Flash (960). As detailed in Table 3, its key achievements include:
  • ASR: Achieves a SOTA CER of 3.82 on Korean benchmarks and a competitive WER of 6.47 on English OpenASR.
  • AST: Achieves SOTA performance on the Fleurs Ko2En task with 19.5 BLEU, outperforming both Whisper and Gemini-2.5-Flash. Furthermore, it demonstrates competitive performance on the Minds14 dataset (28.9 BLEU), comparable to the strong Whisper baseline.
  • SSUM: Delivers competitive scores of 5.40 (English) and 4.12 (Korean), performing on par with several larger SLMs.
  • SQQA: Achieves SOTA performance on the evaluated benchmarks, attaining accuracies of 69.5% on LibriSQA (English) and 64.9% on KorQuAD-speech (Korean)
In summary, FastSLM provides a highly effective and efficient solution for diverse speech-language tasks, proving that a compact speech representation can drive SOTA performance.
4.5 Ablation Study
Effect of Speech Token Compression Ratio on ASR Performance: To determine the optimal speech token compression ratio, we conducted an ablation study that evaluates the trade-off between ASR performance (WER) and computational cost. As illustrated in Fig. 3, a clear relationship emerges. While a high token rate (2.67 tokens/sec) yields the best ASR performance, it incurs a substantial computational cost. In contrast, an overly compressed representation (1.33 tokens/sec) results in a significant degradation of performance.
Figure 3: ASR performance of FastSLM with various speech tokens. (left) LS-clean decoding result, and (right) Voxpopuli decoding result.
Our analysis identifies 1.67 tokens/sec as an optimal operating point that balances these competing factors, achieving strong performance while minimizing computational demands. This decision is supported by a mathematical analysis of the point of diminishing returns, detailed in Appendix J.
Figure 4: Comparison VRAM usage and time-to-first-token (TTFT) according to speech length.
Scaling Limits of Long-Form Speech Input with FastSLM: To provide a practical assessment beyond indirect complexity metrics like parameters and FLOPs (Ma et al., 2018), we empirically evaluate the scaling properties of FastSLM. We measure two key indicators on a single 40GB NVIDIA A100 GPU: VRAM consumption to assess memory efficiency and time-to-first-token (TTFT) to quantify the latency introduced by our multi-layered HFQ-Former.
The results, presented in Fig. 4, highlight significant advantages in scalability. While benchmark models exhibit exponential VRAM growth, FastSLM demonstrates near-linear scaling, successfully processing an 8 hours (28,800 seconds) speech stream using under 30GB of memory. Regarding latency, FastSLM maintains a competitive TTFT. Although its initial latency is comparable to the similarly-sized Voxtral-Mini, it scales far more effectively, showing only a minimal increase as the speech length grows, in contrast to the sharp rise observed in the baseline. These findings confirm that FastSLM architecture enables robust inference on ultra-long-form speech far beyond the capabilities of existing models, especially within constrained environments. Furthermore, its efficiency is highly advantageous for batch processing; the minimal memory footprint per stream allows for significantly larger batch sizes on a single GPU, thereby maximizing throughput for parallelized workloads.
Impact of Hierarchical Modules and Training Strategy: To validate the efficacy of our architectural design and the three-stage training strategy, we conducted an ablation study by selectively removing key components: the Downsampler stage, the hierarchical attention mechanism, and Training Stage 2 (long-form speech adaptation). As summarized in Table 4, removing the Downsampler or hierarchical attention caused a severe performance degradation, with WER on LS-Long (Park et al., 2024) surging to 12.4 and 10.8, respectively. This underscores the structural necessity of our hierarchical compression in handling long sequences. Furthermore, omitting Training Stage 2, which aligns the model with long-context speech via Long-form speech adaptation, resulted in significant regression across tasks—specifically, increasing LS-Long WER from 5.98 to 6.81 and dropping KorQuAD-Speech accuracy from 64.9% to 59.0%. The full FastSLM configuration consistently outperforms all ablated variants, confirming that both the hierarchical architecture and the dedicated adaptation stage are integral to achieving efficient and robust long-form understanding.
Table 4: Comparison of long-form speech adaptation strategy.
Method Dataset LS-Long WER $\downarrow$ KorQuAD-Speech ACC $\uparrow$ SDS-PART6 Score (1-7) $\uparrow$
w/o Downsample Stage $12.4$ $56.7$ $4.12$
w/o Hierarchical Attention $10.8$ $56.9$ $4.92$
w/o Training Stage 2 $6.81$ $59.0$ $5.07$
FastSLM $\mathbf{5.78}$ $\mathbf{64.9}$ $\mathbf{5.40}$
5 Conclusion
In this paper, we introduced FastSLM, a lightweight and efficient SLM engineered to overcome the critical scaling limitations of long-form speech processing. At the core of our approach is the HFQ-Former, a novel module that hierarchically compresses high-frame-rate speech features into compact semantic tokens. Complemented by a cost-effective three-stage training strategy, FastSLM achieves SOTA or competitive performance across diverse benchmarks while reducing feature sequence length by up to 96%. These results demonstrate that extreme token compression is a practical and scalable solution for real-time long-form speech understanding, contributing a critical building block for future multimodal systems. A discussion on current limitations and future directions is provided in Appendix K.
Appendix A Qualitative Analysis of HFQ-Former Attention Map
Fig. 5 shows the cross-attention patterns of HFQ-Former for short-form (top), mid-form (middle), and long-form (bottom) speech across the three hierarchical stages.
For short-form speech, the final queries attend broadly to all stages, indicating that local and mid-term features remain useful when the sequence is short. For mid-form speech, attention begins to shift away from Stage 1 and is increasingly concentrated on Stage 2 and Stage 3, reflecting the need for broader temporal context. For long-form speech, attention becomes strongly dominated by Stage 3, while Stage 1 and Stage 2 receive minimal attention. This shows that the model relies on high-level, compressed representations for long-form speech reasoning.
Overall, as speech length increases, the attention distribution progressively moves from local (Stage 1) to global (Stage 3) features, demonstrating that HFQ-Former adaptively adjusts its focus based on speech length.
Figure 5: Visualization of HFQ-Former attention patterns across speech duration (logarithmic scale). Top: short-form speech ($30\,\mathrm{s}$). Middle: mid-form speech ($60\,\mathrm{s}$). Bottom: long-form speech ($15\,\mathrm{min}$).
Appendix B Additional Analysis of Hierarchical Temporal Modeling
To further substantiate the necessity of hierarchical temporal abstraction in HFQ-Former, we present additional analyses examining how different stages attend to both local and global speech context and compare HFQ-Former against a larger single-stage Q-Former operating under the same token budget. Local temporal cues correspond to short-range phonetic transitions, formant dynamics, and prosodic micro-patterns within 50 to 200 ms. In contrast, global temporal cues capture long-range structure such as discourse flow, topic progression, and speaker turns, which are especially important for long-form reasoning tasks including SSUM and SQQA. A single-stage Q-Former must simultaneously compress thousands of speech frames, forcing local and global cues to compete within a single attention scale, which leads to representational bottlenecks.
HFQ-Former alleviates this limitation through progressive temporal abstraction: early stages attend to fine-grained acoustic details, while deeper stages integrate increasingly broader temporal context. This hierarchical design enables efficient multi-scale modeling without increasing the token budget. Our cross-attention map visualizations (Appendix A) further confirm that attention shifts toward deeper stages as input duration increases, highlighting the importance of multi-stage processing for long-range temporal understanding. The empirical comparison between HFQ-Former and the larger single-stage Q-Former under the same token budget is presented in Table 5.
Table 5: Comparison between a larger single-stage Q-Former and our HFQ-Former under the same token budget.
Method Dataset LS-clean WER $\downarrow$ Voxpopuli WER $\downarrow$ SDS-PART6-Speech Score (1-7) $\uparrow$ KorQuAD-Speech ACC $\uparrow$
Larger Single-stage Q-Former 2.11 7.11 5.04 61.2
HFQ-Former (ours) 2.09 6.55 5.40 64.9
Although the ASR quality (LS-clean) is comparable across the two models, HFQ-Former achieves substantially higher performance on long-form and reasoning-intensive tasks. These results empirically demonstrate that hierarchical temporal decomposition is critical for robust long-form speech understanding.
Appendix C Effect of Hierarchical Staging in HFQ-Former
To directly evaluate the benefit of the hierarchical design in HFQ-Former, we compare three variants using only Stage 1, Stage 1–2, and the full Stage 1–2–3. Table 6 shows that performance consistently improves as more hierarchical stages are included, demonstrating that progressive temporal abstraction is essential for long-form speech processing.
Table 6: Effect of hierarchical downsampling stages on speech understanding performance.
Method Dataset LS-clean WER $\downarrow$ SDS-PART6-Speech Score (1-7) $\uparrow$ KorQuAD-Speech ACC $\uparrow$
Stage 1 2.23 4.12 56.7
Stage 1/2 2.15 4.98 62.2
Stage 1/2/3 2.09 5.40 64.9
Appendix D Prompt Template and Hierarchical Tags for Training
We applied a unified prompt template and task/language control tokens during training to support multiple speech-language tasks:
[…]
To improve task specialization and language awareness, we employed hierarchical task and language tokens, which enabled robust detection and performance across languages and tasks:
$\begin{gathered}\text{Language Token}:<|\text{KO}|><|\text{EN}|>\\ \text{Task Token}:<|\text{ASR}|><|\text{AST}|><|\text{SQQA}|><|\text{SSUM}|>\end{gathered}$
Appendix E Pre-training Dataset
Table 7 is a dataset used for pre-training of FastSLM. The dataset consists of English and Korean, and consists of a total of 10M speech-text pairs.
Table 7: Pre-training dataset details. En denote the English, and Ko denote the Korean.
Dataset Duration (hours) # Samples Speech Language
LibriSpeech 960 281,241 En
TED-LIUM-release3 454 268,263 En
GigaSpeech-L 2,500 2,266,371 En
Voxpopuli 523 182,482 En
SpgiSpeech-M 1,000 385,361 En
Earnings-22 105 52,006 En
AMI 78 108,502 En
Common Voice 15 2,532 1,070,066 En
AI-HUB ASR-En 1,000 1,020,265 En
AI-HUB ASR-Ko 7,812 4,557,512 Ko
Total 16,964 10,212,348 -
Appendix F Prompt for GPT-4 as a Judge on Speech Benchmarks
The following is the exact prompt template used for evaluating SSUM output via an LLM-as-a-Judge.
Listing 1: LLM-as-a-Judge Prompt for SSUM Evaluation
You are a skilled evaluator for summaries generated based on user-provided instructions.
Your task is to rate how well the summary follows the users instructions on a 1-7 scale.
Scoring Rubric:
- **7 (Excellent):** Fully follows all instructions. Accurate, fluent, and coherent with the correct level of detail and structure.
- **6 (Good):** Almost perfect, with very minor issues that do not affect usability (e.g., tiny structural deviation, trivial omission).
- **5 (Mostly Correct):** Fulfills the main instruction but has noticeable issues (e.g., includes some unimportant extras, misses a few details).
- **4 (Acceptable):** Adheres to the instruction partially but has significant issues like inconsistencies or irrelevant content.
- **3 (Poor):** Minimally adheres to the instruction, missing most required details or containing significant irrelevant/hallucinated content.
- **2 (Very Poor):** Fails to follow the core instruction. Mostly irrelevant, fabricated, or ignores requested structure/tone.
- **1 (Fails):** Completely fails to follow instructions.
Input:
- **User Instruction:** {USER_INSTRUCTION}
- **Reference (gold):** {REFERENCE_ANSWER}
- **Model Summarization:** {SUMMARY_TO_EVALUATE}
Notes:
- It helps to read the Summary first, then compare with the Reference and Instruction.
- If the summary is missing or empty, return N/A as the score.
Output:
Note: Use the following JSON format for easy downstream consumption.
explanation: Brief reasoning for the score based on the rubric.,
score: <Float, 1-7>
Appendix G Generation and Decoding Configuration
We use the following decoding hyperparameters for all LLM-based generation tasks.
Table 8: Decoding configuration used for LLM-based generation in FastSLM.
Parameter Value
Decoding Strategy Sampling
Temperature 0.2
Top-p 0.95
Top-k 20
Repetition Penalty 1.0
Appendix H Model and Training Parameters
The model implementation details and the training setup for each stage are presented in Table 9 and Table 10.
Table 9: Model configuration for FastSLM.
Module Component Configuration
Encoder Backbone Parameters Hidden Size Context Length Whisper-large-v3 635M 1280 1500
Adapter Backbone Parameters Hidden Size Queries per Stage Compressed Speech Token Downsampling Factors HFQ-Former 56M 1280 80 50 2
LLM Backbone Parameters Hidden Size Context Length Qwen3-4B 4.06B 2560 4096
LoRA Rank ($r$) Alpha ($\alpha$) Scaling Factor LoRA Target Modules 16 64 4 q/k/v_proj, gate/up/down_proj
Table 10: Training settings across stages
Setting Stage1 Stage2 Stage3
Learning Rate 1e-4 5e-5 5e-5
Learning Rate Scheduler Linear Decay
Weight Decay 0 1e-4 1e-4
Epoch 1 1 2
Data Type BF16
DeepSpeed Stage Zero2
Appendix I Details of ASR Benchmark Results
Table 11 presents a detailed comparison of FastSLM and SOTA models across multiple ASR benchmarks. We report WER for English datasets and CER for Korean datasets. The results demonstrate that FastSLM achieves competitive performance while using significantly fewer speech tokens per second.
Table 11: Comparison of WER between FastSLM and state-of-the-art (SOTA) models. This results representation ASR Benchmark Dataset WER and CER.
Dataset Sub-Category Metric FastSLM 4.8B Qwen2-Audio 8B Phi4-Multimodal 5.8B Whisper 1.5B Voxtral-mini 4.7B Gemini-2.5- Flash
OpenASR AMI WER $10.8$ $15.2$ $11.7$ $16.0$ $16.3$ $21.6$
Earnings22 WER $10.7$ $14.1$ $10.2$ $11.3$ $10.7$ $13.1$
GigaSpeech WER $10.7$ $10.3$ $9.78$ $10.0$ $10.2$ $10.7$
SpgiSpeech WER $2.33$ $3.00$ $3.13$ $2.01$ $2.37$ $3.82$
TEDLIUM WER $3.97$ $4.05$ $2.90$ $3.91$ $3.68$ $3.01$
LS-clean WER $2.09$ $1.74$ $1.68$ $2.94$ $1.88$ $2.49$
LS-other WER $4.67$ $4.03$ $3.83$ $3.86$ $4.10$ $5.84$
Voxpopuli WER $6.55$ $7.05$ $5.91$ $9.54$ $7.14$ $7.89$
Fleurs En WER $5.26$ $5.27$ $3.38$ $4.10$ $3.77$ $6.20$
Ko CER $2.79$ N/A N/A $5.32$ N/A 3.00
Common Voice 15 En WER $10.9$ $8.68$ $7.61$ $9.30$ $10.2$ $11.2$
Ko CER $4.55$ N/A N/A $5.74$ N/A $6.09$
Appendix J Mathematical Justification for Speech Token Ratio
To provide a more mathematical justification for our choice of 1.67 speech tokens/sec, we analyze the marginal gain in performance versus the marginal increase in cost, identifying the point of diminishing returns. We define a simple Efficiency Score to formalize this trade-off:
$\text{Efficiency Score}=\frac{\Delta\text{Performance (WER Reduction)}}{\Delta\text{Cost (FLOPS Increase)}}$
We apply this metric to the data from the LS-clean chart (Fig. 3). The results, summarized in Table 12, clearly show a sharp drop in efficiency after the 1.67 tokens/sec interval.
Table 12: Efficiency Score calculation for different token rate intervals on the LS-clean dataset. A higher score indicates greater efficiency.
[…]
$1.33\rightarrow 1.67$ $\approx 0.3\%$ $\approx 0.3$ T $\approx\mathbf{1.00}$
$1.67\rightarrow 2.00$ $\approx 0.1\%$ $\approx 1.3$ T $\approx\mathbf{0.08}$
The analysis in Table 12 quantitatively demonstrates that the interval beyond 1.67 tokens/sec marks a stark point of diminishing returns. Although adding more tokens continues to slightly lower the WER, the computational cost required for each marginal improvement becomes disproportionately high. Therefore, 1.67 tokens/sec is the most efficient configuration, maximizing the performance gain before the cost-benefit ratio sharply declines.
Appendix K Limitation
While FastSLM demonstrates a significant step forward in creating computationally efficient and scalable SLMs, this work has several limitations that warrant consideration and offer avenues for future research.
First, our approach to enabling Korean-language capabilities relies heavily on instruction-tuning datasets generated with a TTS engine, which introduces a potential synthetic-to-real domain gap. Synthetic speech generally lacks natural prosody, spontaneous disfluencies (e.g., hesitations, restarts), background noise, and speaker variability that occur in real conversational speech. As a result, although FastSLM shows strong performance on our synthetic Korean evaluation sets, its ability to generalize to authentic, real-world Korean speech remains unverified and may be significantly lower. Moreover, because evaluation is largely performed on data drawn from the same distribution as the training data, current metrics may overestimate real-world performance.
Second, FastSLM is evaluated exclusively on speech tasks. While this focus enables high efficiency and strong performance for long-form speech, the model has not been assessed on broader auditory scenes (e.g., environmental sounds or music). Extending FastSLM toward general audio understanding remains an important direction for future work.
To address this limitation, we plan to extend training and evaluation to audio understanding tasks, enabling a more comprehensive assessment of the FastSLM ability to preserve detailed acoustic information.
Appendix L Use of AI Writing Assistance
LLMs were used solely to improve the language and clarity of this manuscript. All technical content was authored by the human authors, and all AI-assisted edits were reviewed and approved by the authors.